# ASR-enhanced Multimodal Representation Learning for Cross-Domain Product Retrieval

### v1

@@ IV-B Comparison with Baseline Methods
Baselines. Based on our analysis of current methods for CdPR, c.f. Table II, COPE is a natural baseline to AMPere. Moreover, we compare with CN-CLIP [36]. This Chinese variant of CLIP is a widely accepted baseline for both generic text-to-video retrieval [37] and CdPR [3, 13, 2]. Per video, we obtain its CN-CLIP feature by mean pooling over the 8 frame-level features. The multimodal feature is obtained by averaging the video-level feature and the CN-CLIP feature of the associated ASR text. To check if LLM-summarized text is also beneficial for CN-CLIP, we implement three variants:
  • CN-CLIP-m0: multimodal feature obtained by the raw ASR text under zero-shot setting.
  • CN-CLIP-m1: multimodal feature obtained by the LLM-summarized ASR text under zero-shot setting.
@@
We further compare with QCD [2] under two settings:
@@
Results. The performance of the different models is shown in Table IV. Given COPE as the baseline, AMPere improves mR1 from 56.4 to 62.1, i.e. a relative improvement of 10.5%. The attention maps in Fig. 3 reveal that AMPere is more focused on product regions. Compared to COPE, AMPere has a smaller intra-product distance (0.53 vs. 0.46) and a larger inter-product distance (0.96 vs. 0.99), see Fig. 4, showing that the latter yields a more discriminative space for cross-domain product representation. See Fig. 5 for qualitative results. The superiority of AMPere over COPE is thus verified.
TABLE IV: Comparison with SOTA for cross-domain product retrieval.
. Model P2S P2L S2P S2L L2P L2S mR1 R1 R5 R10 R1 R5 R10 R1 R5 R10 R1 R5 R10 R1 R5 R10 R1 R5 R10 Visual-only: COPE 82.6 94.9 97.5 54.1 71.1 77.1 65.2 76.6 82.0 46.0 63.6 70.6 42.3 56.5 63.7 48.3 67.2 74.7 56.4 CN-CLIP 58.8 78.6 85.5 32.6 46.1 52.9 35.7 53.1 60.5 26.0 39.0 50.0 18.9 33.4 40.5 28.4 46.6 54.5 33.4 (40.1% $\downarrow$) Multimodal: CN-CLIP-m0 55.5 82.4 90.1 38.3 62.7 71.3 19.7 34.0 41.2 22.8 39.6 48.7 9.3 17.1 22.0 15.4 33.4 43.4 26.8 (52.5% $\downarrow$) CN-CLIP-m1 63.2 85.7 92.4 48.5 70.0 78.3 34.6 52.3 59.1 30.1 46.6 54.1 25.5 42.1 49.9 31.8 52.4 61.6 39.0 (30.9% $\downarrow$) CN-CLIP-m1-finetune 76.1 92.3 96.0 52.5 71.6 79.0 64.2 79.0 64.6 45.5 64.3 72.3 42.1 58.8 66.7 44.5 65.9 73.7 54.2 (3.9% $\downarrow$) QCD 69.3 89.2 94.5 52.3 74.4 82.5 63.7 79.2 84.4 47.3 67.7 75.5 42.1 59.2 66.3 43.0 65.3 73.5 52.9 (6.2% $\downarrow$) QCD+ 75.5 91.9 96.3 60.7 78.1 84.4 69.1 80.5 86.1 53.4 69.2 76.3 47.4 63.0 70.6 49.6 67.2 75.5 59.3 (5.1% $\uparrow$) mPLUG-Video-0 0.1 0.5 0.8 0.4 1.0 1.2 0.2 0.5 0.6 0.4 1.0 1.1 0.1 0.4 0.5 0.1 0.4 0.6 0.2 (99.6% $\downarrow$) mPLUG-Video-lb 50.7 76.7 86.1 38.3 63.2 73.1 40.7 62.5 71.6 29.8 51.1 62.2 26.2 46.9 58.4 27.5 50.3 60.9 35.5 (37.1% $\downarrow$) Proposed AMPere 84.5 95.9 98.3 65.0 81.4 86.5 71.2 84.3 88.6 54.7 71.3 77.5 47.8 64.6 72.2 50.4 68.9 76.2 62.3 (10.5% $\uparrow$)
@@ IV-C Understanding AMPere
As shown in Table V, LLM performs the best in both scenarios. We observe that the LLM-summarized text is richer than its UIE counterpart, see Fig. 3. This is also supported by the fact that the LLM-summarized text is on average much longer, see Table III. Given the uncontrolled nature of the ASR text, the LLM-based text summarizer may yield no result. However, the chance of no result is much lower than its UIE counterpart, see Table VI.
@@
Considering the multimodal scenario, the gain of adding the raw ASR (0.7%) or the UIE-summarized text (1.6%) compared to the visual-only solution, i.e. COPE, is marginal. By contrast, using the LLM-summarized text obtains a relative improvement of 5.8%, see the last row of Table V. The result verifies the necessity of ASR text processing and the effectiveness of our ASR text summary.
TABLE V: Evaluating different usages of ASR text.
@@
Evaluation of the three text summarizers. In order to directly evaluate the output of the three ASR text summarizers, we build a ground-truthed dataset as follows. From the test set, we randomly selected 200 raw ASR text samples from the short-video domain and the live-stream domain, respectively. Recall that the text summarizers are designed to extract two product-specific fields, i.e. product-name and product-features. For product name extraction, we manually identify the product name by jointly checking each text sample and its corresponding video. Note that some videos are merely associated with background music or casual chat. Eventually, we successfully identified product names for 133 text samples from the short-video domain and 163 text samples from the live-stream domain. As shown in Table VI, our LLM-based text summarizer, with lower no-output rate and higher accuracy, is clearly better than its KeyBERT and UIE counterparts.
TABLE VI: Evaluating quality of product name extraction.
@@
No-output rate Accuracy No-output rate Accuracy No-output rate Accuracy
@@
UIE 40.6% 48.1% 47.9% 36.2% 44.6% 41.6%
LLM 4.5% 93.2% 5.5% 91.4% 5.1% 92.2%
In addition, the performance of product features (attributes) extraction is reported in Tab. VII. As a specific product is typically associated with multiple attributes, making it essentially a multi-label problem, we report both recall and accuracy. Our LLM-based text summarize also clearly outperforms KeyBERT and UIE in product features (attributes) extraction.
TABLE VII:
Evaluating quality of product attribute extraction
@@
Ablations of other LLMs. We investigate other LLMs by directly using the AMPere model trained with Baichuan2-13B-Chat-4bits cleaned text to see how our model works when replacing the current LLM in summarizer by two smaller alternatives at the inference time, i.e. Baichuan2-7B-Chat-4bits and Qwen2-1.5B-Instruct [^8] https://github.com/QwenLM/Qwen2. As Tab. VIII shows, at a relatively small loss of mR1 (60.2 $\rightarrow$ 58.4, 3.0%), GPU footprint is reduced by 12.0% (13.7GB $\rightarrow$ 12.0GB).
TABLE VIII: Ablations of other LLMs.
@@
As shown in Table IX, our current choice is on par with CoA, yet uses only half the number of parameters.
TABLE IX:
@@
As Table X shows, parameter sharing is better. The same conclusion is obtained when training the unimodal models, either visual or textual. Our interpretation is that when the parameters are not shared, the three branches tend to produce domain-specific embeddings, thus making cross-domain embedding alignment more challenging. So, in contrast to [3], we suggest parameter sharing.
TABLE X: Evaluating the parameter sharing strategy.
@@
Contribution per domain. To reveal the influence of the individual domains on MmPRL, we separately exclude each domain from training. As Table XI shows, the short-video domain is the most important, followed by the product-page domain and the live-stream domain. With manual inspection, we observe that in the ROPE dataset, the short-video domain has a considerable amount of live-stream style videos. Whether this phenomenon stems from biases in the dataset gathering process or properties inherent to the E-commerce platform where the dataset was sourced requires further investigation.
TABLE XI: Influence of the individual domains when excluded from training. The short-video domain is the most influential. The live-stream domain is the least important.
@@
(a) S2P (product: electric shaver)
(b) L2P (product: shower head)
(c) P2S (product: triangle clamp)
(d) P2L (product: wedding balloon)
(e) L2S (product: pineapple)
(f) S2L (product: scalable football)
Figure 5: Hand-picked qualitative results of the six CdPR tasks. Per figure, from top to bottom, are top-5 results retrieved by COPE, AMPere with raw ASR, and AMPere, respectively. Text associated with the visual samples is not shown for simplicity. Correct results are shown with green borders. Best viewed on screen.
V Conclusions and Remarks
@@
Limitations of this study. As we have noted, due to the uncontrolled nature of the ASR text, the LLM-based text summarizer yielded no result for the samples. In the current work, we simply use an empty string as a hotfix. How to better handle the modality missing issue warrants further research.

### v2

@@ IV-B Comparison with Baseline Methods
Baselines. Based on our analysis of current methods for CdPR, c.f. Table II, COPE is a natural baseline to AMPere. Moreover, we compare with CN-CLIP [42]. This Chinese variant of CLIP is a widely accepted baseline for both generic text-to-video retrieval [43] and CdPR [3, 15, 2]. Per video, we obtain its CN-CLIP feature by mean pooling over the 8 frame-level features. To check if raw ASR text or LLM-summarized text is also beneficial for CN-CLIP, we implement three variants:
  • CN-CLIP-m0: Multimodal feature is obtained by averaging the video-level feature and the CN-CLIP feature of associated raw ASR text, under zero-shot setting.
  • CN-CLIP-m1: Similar to CN-CLIP-m0, yet using LLM-summarized ASR text instead of raw ASR text.
@@
We further compare with QCD [2], another multimodal method, under two settings:
@@
Note that except for COPE and CN-CLIP, all the other baselines have utilized ASR text (or its clean version) for multimodal representation.
Results. The performance of the different models is shown in Table IV. Given COPE as the baseline, AMPere improves mR1 from 56.4 to 62.1, i.e. a relative improvement of 10.5%. Our model is also the best in terms of MRR and NDCG10, see Table V and VI. The attention maps in Fig. 3 reveal that AMPere is more focused on product regions. Compared to COPE, AMPere has a smaller intra-product distance (0.53 vs. 0.46) and a larger inter-product distance (0.96 vs. 0.99), see Fig. 4, showing that the latter yields a more discriminative space for cross-domain product representation. See Fig. 5 and Fig. 6 for qualitative results. The superiority of AMPere over COPE is thus verified.
TABLE IV: Comparison with SOTA.
Model P2S P2L S2P S2L L2P L2S mR1
R1 R5 R10 R1 R5 R10 R1 R5 R10 R1 R5 R10 R1 R5 R10 R1 R5 R10
Visual-only:
COPE 82.6 94.9 97.5 54.1 71.1 77.1 65.2 76.6 82.0 46.0 63.6 70.6 42.3 56.5 63.7 48.3 67.2 74.7 56.4
Cross-modal:
CN-CLIP 58.8 78.6 85.5 32.6 46.1 52.9 35.7 53.1 60.5 26.0 39.0 50.0 18.9 33.4 40.5 28.4 46.6 54.5 33.4 (40.1% $\downarrow$)
Multimodal:
CN-CLIP-m0 55.5 82.4 90.1 38.3 62.7 71.3 19.7 34.0 41.2 22.8 39.6 48.7 9.3 17.1 22.0 15.4 33.4 43.4 26.8 (52.5% $\downarrow$)
CN-CLIP-m1 63.2 85.7 92.4 48.5 70.0 78.3 34.6 52.3 59.1 30.1 46.6 54.1 25.5 42.1 49.9 31.8 52.4 61.6 39.0 (30.9% $\downarrow$)
CN-CLIP-m1-finetune 76.1 92.3 96.0 52.5 71.6 79.0 64.2 79.0 64.6 45.5 64.3 72.3 42.1 58.8 66.7 44.5 65.9 73.7 54.2 (3.9% $\downarrow$)
QCD 69.3 89.2 94.5 52.3 74.4 82.5 63.7 79.2 84.4 47.3 67.7 75.5 42.1 59.2 66.3 43.0 65.3 73.5 52.9 (6.2% $\downarrow$)
QCD+ 75.5 91.9 96.3 60.7 78.1 84.4 69.1 80.5 86.1 53.4 69.2 76.3 47.4 63.0 70.6 49.6 67.2 75.5 59.3 (5.1% $\uparrow$)
mPLUG-Video-0 0.1 0.5 0.8 0.4 1.0 1.2 0.2 0.5 0.6 0.4 1.0 1.1 0.1 0.4 0.5 0.1 0.4 0.6 0.2 (99.6% $\downarrow$)
mPLUG-Video-lb 50.7 76.7 86.1 38.3 63.2 73.1 40.7 62.5 71.6 29.8 51.1 62.2 26.2 46.9 58.4 27.5 50.3 60.9 35.5 (37.1% $\downarrow$)
Proposed AMPere 84.5 95.9 98.3 65.0 81.4 86.5 71.2 84.3 88.6 54.7 71.3 77.5 47.8 64.6 72.2 50.4 68.9 76.2 62.3 (10.5% $\uparrow$)
TABLE V:
Comparison with SOTA. Metric: MRR.
Model P2S P2L S2P S2L L2P L2S Mean
COPE 0.892 0.639 0.761 0.596 0.503 0.580 0.662
CN-CLIP 0.678 0.397 0.440 0.326 0.262 0.372 0.412
CN-CLIP-m0 0.673 0.494 0.270 0.313 0.137 0.243 0.355
CN-CLIP-m1 0.731 0.585 0.430 0.383 0.336 0.415 0.480
CN-CLIP-m1-finetune 0.831 0.615 0.710 0.544 0.504 0.545 0.625
QCD 0.782 0.599 0.679 0.569 0.496 0.510 0.606
QCD+ 0.830 0.664 0.763 0.617 0.541 0.585 0.667
mPLUG-Video-0 0.005 0.005 0.005 0.006 0.003 0.005 0.005
mPLUG-Video-lb 0.624 0.500 0.508 0.402 0.364 0.385 0.464
AMPere 0.895 0.726 0.767 0.628 0.558 0.595 0.695
TABLE VI: Comparison with SOTA. Metric: NDCG10.
Model P2S P2L S2P S2L L2P L2S Mean
COPE 0.763 0.503 0.680 0.470 0.399 0.437 0.542
CN-CLIP 0.398 0.285 0.319 0.233 0.167 0.208 0.268
CN-CLIP-m0 0.428 0.322 0.168 0.190 0.074 0.113 0.216
CN-CLIP-m1 0.516 0.419 0.318 0.258 0.231 0.246 0.332
CN-CLIP-m1-finetune 0.655 0.474 0.607 0.419 0.396 0.384 0.489
QCD 0.577 0.486 0.632 0.427 0.365 0.328 0.469
QCD+ 0.635 0.565 0.687 0.488 0.457 0.446 0.546
mPLUG-Video-0 0.003 0.001 0.002 0.002 0.001 0.001 0.002
mPLUG-Video-lb 0.389 0.328 0.362 0.257 0.249 0.221 0.301
AMPere 0.778 0.596 0.699 0.500 0.475 0.452 0.583
@@ IV-C Understanding AMPere
As shown in Table VII, LLM performs the best in both scenarios. We observe that the LLM-summarized text is richer than its UIE counterpart, see Fig. 3. This is also supported by the fact that the LLM-summarized text is on average much longer, see Table III. Given the uncontrolled nature of the ASR text, the LLM-based text summarizer may yield no result. However, the chance of no result is much lower than its UIE counterpart, see Table VIII.
@@
Considering the multimodal scenario, the gain of adding the raw ASR (0.7%) or the UIE-summarized text (1.6%) compared to the visual-only solution, i.e. COPE, is marginal. By contrast, using the LLM-summarized text obtains a relative improvement of 5.8%, see the last row of Table VII. The result verifies the necessity of ASR text processing and the effectiveness of our ASR text summary. The superior performance of AMPere against the multimodal baseline of “Visual + Raw ASR” (60.2 vs. 57.3 in mR1) allows us to attribute the improvement to the advantages of AMPere rather than the inclusion of ASR text.
TABLE VII: Evaluating different usages of ASR text.
@@
Evaluation of the three text summarizers. In order to directly evaluate the output of the three ASR text summarizers, we build a ground-truthed dataset as follows. From the test set, we randomly selected 200 raw ASR text samples from the short-video domain and the live-stream domain, respectively. Recall that the text summarizers are designed to extract two product-specific fields, i.e. product-name and product-features. For product name extraction, we manually identify the product name by jointly checking each text sample and its corresponding video. Note that some videos are merely associated with background music or casual chat. Eventually, we successfully identified product names for 133 text samples from the short-video domain and 163 text samples from the live-stream domain. As Tab. VIII shows, our LLM-based text summarizer is clearly better than its KeyBERT and UIE counterparts.
TABLE VIII: Accuracy of different ASR text summarizers for product-name extraction.
@@
UIE 48.1% 36.2% 41.6%
LLM 93.2% 91.4% 92.2%
The performance of product features (attributes) extraction is reported in Tab. IX. As a specific product is typically associated with multiple attributes, making it essentially a multi-label problem, we report both recall and accuracy. Our LLM-based text summarize also clearly outperforms KeyBERT and UIE in product features (attributes) extraction.
TABLE IX: Evaluating product-attribute extraction.
@@
Ablations of other LLMs. We investigate other LLMs by directly using the AMPere model trained with Baichuan2-13B-Chat-4bits cleaned text to see how our model works when replacing the current LLM in summarizer by two smaller alternatives at the inference time, i.e. Baichuan2-7B-Chat-4bits and Qwen2-1.5B-Instruct [^8] https://github.com/QwenLM/Qwen2. As Tab. X shows, at a relatively small loss of mR1 (60.2 $\rightarrow$ 58.4, 3.0%), GPU footprint is reduced by 12.0% (13.7GB $\rightarrow$ 12.0GB).
TABLE X: Ablations of other LLMs.
@@
As shown in Table XI, our current choice is on par with CoA, yet uses only half the number of parameters.
TABLE XI:
@@
As Table XII shows, parameter sharing is better. The same conclusion is obtained when training the unimodal models, either visual or textual. Our interpretation is that when the parameters are not shared, the three branches tend to produce domain-specific embeddings, thus making cross-domain embedding alignment more challenging. So, in contrast to [3], we suggest parameter sharing.
TABLE XII: Evaluating the parameter sharing strategy.
@@
Contribution per domain. To reveal the influence of the individual domains on MmPRL, we separately exclude each domain from training. As Table XIII shows, the short-video domain is the most important, followed by the product-page domain and the live-stream domain. With manual inspection, we observe that in the ROPE dataset, the short-video domain has a considerable amount of live-stream style videos. Whether this phenomenon stems from biases in the dataset gathering process or properties inherent to the E-commerce platform where the dataset was sourced requires further investigation.
TABLE XIII: Influence of the individual domains when excluded from training. The short-video domain is the most influential. The live-stream domain is the least important.
@@
Robustness analysis. In order to reveal how AMPere performs under varying ASR data quality, per ASR text we calculate its signal level as the ratio of the LLM-summarized text length to the raw ASR text length. A bigger ratio indicates a larger proportion of product-related information within the raw text, and thus a stronger signal level. Accordingly, we group test queries according to the signal level of their associated ASR text. As shown in Table XIV, AMPere consistently outperforms the baselines at varied ASR signal levels. Even in the extreme case of ZERO signal, i.e. LLM no-output, AMPere is also better than COPE (38.1 vs 36.2 in R1). We attribute this result to the fact that with multimodal learning, the visual encoder of AMPere is stronger than its counterpart in COPE, see Fig. 3. As such, even the test queries lack meaningful ASR text input, using the multimodal model remains beneficial.
TABLE XIV: Performance under varying ASR data quality. Query domain: L / Retrieval domain: P. Metric: R1.
ASR Signal Level Percentage Visual only Visual + Raw ASR AMPere
0 (LLM No-output) 3.4% 36.2 33.4 38.1
$(0,0.05)$ 10.4% 35.5 37.3 39.2
$[0.05,0.1)$ 48.1% 43.7 43.3 48.2
$[0.1,0.15)$ 23.1% 43.6 44.3 47.3
$[0.15,0.2)$ 8.2% 37.6 40.6 46.5
$[0.2,0.25)$ 3.7% 39.7 39.1 42.8
$>=0.25$ 3.1% 38.9 40.7 41.3
Failure case analysis. By manually inspecting the retrieval results of the current worst-performing task, i.e. L2P, we observe three common types of errors as follows:
$\bullet$ Failure Type-I: Multi-product presence. As shown in Fig. 5, when multiple products co-present in a query video, e.g. a toilet brush as the current product and a toilet as a surrounding item, the model may incorrectly retrieve product pages of the surrounding item. To resolve such an ambiguity, product localization is required.
$\bullet$ Failure Type-II: Insufficient multimodal fusion. As illustrated in Fig. 5, although our LLM based text summarizer has successfully extracted product names, e.g. solid bracelet, the result that our model returns as the first hit a photo of a hand with a ring suggests insufficient fusion of the visual and textual inputs. To fully unlock the value of the ASR text, adaptive multimodal fusion requires further research.
$\bullet$ Failure Type-III: Inaccurate text. This type of failures occurs when the LLM-summarized text, e.g. “power bank”, is not sufficiently accurate to describe the product of “pocket-portable power bank” shown in the live stream, see Fig. 5. How to incorporate the visual context into ASR text summarization for more accurate and fine-grained textual information extraction deserves future exploration.
(a) Failure Type-I: Multi-product presence
(b) Failure Type-II: Insufficient multimodal fusion
(c) Failure Type-III: Inaccurate text
Figure 5: Failure cases. Top-1 item is shown. Task: L2P.
Efficiency analysis. The computational cost of training and inference is given in Table XV. Due to the inclusion of the text branch and the use of LLM, AMPere is relatively slower than COPE. It is worth pointing out that in a practical E-commerce system, instances of specific products are known in advance. Hence, their multimodal embeddings are offline computed and stored for CdPR on the fly. In fact, AMPere has already been deployed in Kuaishou E-commerce for real-time product retrieval and recommendation.
TABLE XV: Efficiency analysis. Training 16$\times$×V100. Inference on a V100.
Model Training per epoch Inference per (P,S,L) triplet
GPU footprint Time cost GPU footprint LLM Embedding
COPE 20.0$\times$16 GB 22.2 min 4.6 GB n.a. 49.7 ms
AMPere 22.1$\times$16 GB 31.4 min 5.1 GB 4.4 s 61.7 ms
(a) S2P (electric shaver)
(b) L2P (shower head)
(c) P2S (triangle clamp)
(d) L2S (pineapple)
(e) P2L (wedding balloon)
(f) S2L (scalable football)
Figure 6: Hand-picked qualitative results. Per sub-figure, from top to bottom, are top-5 results retrieved by COPE, AMPere with raw ASR, and AMPere, respectively. Correct results are shown in green. Best viewed on screen.
Potential ethical issues of using ASR data. Although E-commerce network anchors are trained to follow professional ethics and codes of conduct, we cannot rule out the possibility that ASR data might occasionally include personally identifiable information, other sensitive, controversial or even offensive content. We recommend the local deployment of the LLM-based ASR text summarizer as in our current implementation. However, when a proprietary, cloud-based LLM is preferred, an ethics-oriented pre-filtering on the ASR text is necessary before submitting the text to the remote LLM.
V Conclusions